Solve the system by substitution, elimination, or graphing:
4x-3y=2
3x+3y=-9

Answer by CharlesG2(834)

Solve the system by substitution, elimination, or graphing:
4x-3y=2
3x+3y=-9

below is a method of elimination then substitution

the y is eliminated, the x is solved for, then that value of x is substituted
to solve for y

4x - 3y = 2
3x + 3y = -9

add the 2 equations
7x = -7
solve for x
x = -1

plug in -1 into the first equation
4(-1) - 3y = 2
solve for y
-4 - 3y = 2
-3y = 2 + 4
-3y = 6
y = -2

check
4(-1) - 3(-2) = -4 + 6 = 2, yes
3(-1) + 3(-2) = -3 - 6 = -9, yes